find the slope and y intercept of the line and use the slope and y intercept to sketch a graph of the line.
y = 5/3x - 4

Answer by ewatrrr(10682)

 
Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b where m is the slope and b the y-intercept.
find the slope and y intercept of the line and use the slope and y intercept to sketch a graph of the line.
y = (5/3)x - 4 m = 5/3 and y-intercept is (0,-4)
Note: as y increases by 5, x increases by 3. Ex. (3,1) on this Line